OpenStudy (anonymous):

I need help I have a big test tomorrow and I forgot how to do things... sec 39˚ =

OpenStudy (phi):

this is a memory thing: secant(x) = sec(x) = 1/cosine(x) cosecant(x) = csc(x) = 1/sin(x) cot(x) = 1/tan(x) you then use a calculator to find the actual value.

OpenStudy (anonymous):

Okay thank you. :)

OpenStudy (anonymous):

Is the answer 1.29...?

OpenStudy (phi):

yes, if you round to the nearest tenth.